OpenStudy (anonymous):

Suppose the radius of a circle is 4. What is its circumference?

OpenStudy (theprofessor35):

Circumference = π × diameter = 2 × π × radius. T

OpenStudy (aryavegapunk):

Formula for Circumference is 2pi*r where r is the radius

OpenStudy (theprofessor35):

2 * 3.14 * 4 = ?

OpenStudy (theprofessor35):

pi = about 3.14

OpenStudy (theprofessor35):

2*3.14* 4 = 25.12

OpenStudy (theprofessor35):

Now, we know that our circumference is about 25.12

OpenStudy (theprofessor35):

Hope I helped! ^_^